OpenStudy (anonymous):

Which of the following represents x =1/3 y written in general form?

OpenStudy (abbles):

Isolate the y to get it into y-intercept form. Do you know how to do this?

OpenStudy (anonymous):

yea i do

OpenStudy (mathmale):

First: Do you mean x = 1 / (3y) or x = (1/3)y? Please use parentheses freely so that your intent is clear.

OpenStudy (anonymous):

(1/3)y

OpenStudy (mathmale):

OK: Solve x= (1/3)y for y. One way would be to identify the LCD and multiply both sides of this equation by it. There are other approaches. Mind getting started?